OpenStudy (therealmeeeee):

Solve for x. 2 + x < 7 x > 6 x > 5 x < 5 x < 6 @Isaiah.Feynman

OpenStudy (isaiah.feynman):

Subtract 2 from both sides.

OpenStudy (bibby):

I dunno but I have a spare shirt lying around. Wanna trade it for some swag? I'm sure you have more than you need

OpenStudy (the_fizicx99):

2 + x < 7 -2 -2 Can you solve? lol bib

OpenStudy (therealmeeeee):

subtract wat from wat

OpenStudy (isaiah.feynman):

Subtract 2 from both sides of the inequality sign!

OpenStudy (therealmeeeee):

5

OpenStudy (isaiah.feynman):

Yes, 5. It is greater or less than x?

OpenStudy (therealmeeeee):

greater than

OpenStudy (therealmeeeee):

so C

OpenStudy (isaiah.feynman):

Good!

OpenStudy (therealmeeeee):

is it C

OpenStudy (isaiah.feynman):

Yes.

OpenStudy (therealmeeeee):

Which value for x makes the sentence true? 10x - 6 = 5x + 4 2 4 8 -4

OpenStudy (isaiah.feynman):

See, how about you substitute those options one by one to see if they make the expression true.

OpenStudy (therealmeeeee):

wat

OpenStudy (therealmeeeee):

i dont understand

OpenStudy (isaiah.feynman):

Substitute the choices for x and see if the expression on the left equals the expression the right.

OpenStudy (therealmeeeee):

i got B

OpenStudy (isaiah.feynman):

Okay lets see if its true..\[10(4)-6=5(4)+4\] \[40-6=20+4\]\[34=24\]Oops doesn't look right

OpenStudy (therealmeeeee):

oh

OpenStudy (therealmeeeee):

ia it C

OpenStudy (isaiah.feynman):

The aim here is make sure either side of the equal sign have THE SAME value.

OpenStudy (therealmeeeee):

omg its 2

OpenStudy (therealmeeeee):

:)(:

OpenStudy (isaiah.feynman):

How did you know?

OpenStudy (therealmeeeee):

i just did wat u did

OpenStudy (isaiah.feynman):

Smart.

OpenStudy (therealmeeeee):

hahah

OpenStudy (therealmeeeee):

wait wayttttttttttttt

OpenStudy (anonymous):

its 5!

OpenStudy (anonymous):

Specifically C